# Language Models use Lookbacks to Track Beliefs

### v1

@@ Abstract
How do language models (LMs) represent characters’ beliefs, especially when those beliefs may differ from reality? This question lies at the heart of understanding the Theory of Mind (ToM) capabilities of LMs. We analyze Llama-3-70B-Instruct’s ability to reason about characters’ beliefs using causal mediation and abstraction. We construct a dataset that consists of simple stories where two characters each separately change the state of two objects, potentially unaware of each other’s actions. Our investigation uncovered a pervasive algorithmic pattern that we call a lookback mechanism, which enables the LM to recall important information when it becomes necessary. The LM binds each character-object-state triple together by co-locating reference information about them, represented as their Ordering IDs (OIs) in low rank subspaces of the state token’s residual stream. When asked about a character’s beliefs regarding the state of an object, the binding lookback retrieves the corresponding state OI and then an answer lookback retrieves the state token. When we introduce text specifying that one character is (not) visible to the other, we find that the LM first generates a visibility ID encoding the relation between the observing and the observed character OIs. In a visibility lookback, this ID is used to retrieve information about the observed character and update the observing character’s beliefs. Our work provides insights into the LM’s belief tracking mechanisms, taking a step toward reverse-engineering ToM reasoning in LMs.
@@ 1 Introduction
The ability to infer the mental states of others—known as Theory of Mind (ToM)—is an essential aspect of social and collective intelligence (Premack & Woodruff, 1978; Riedl et al., 2021). Recent studies have established that language models (LMs) can solve some tasks requiring ToM reasoning (Street et al., 2024; Strachan et al., 2024a; Kosinski, 2024), while others have highlighted shortcomings (Sclar et al., 2025; Shapira et al., 2024; Kim et al., 2023a, inter alia). Nonetheless, most existing work relies on behavioral evaluations, which do not shed light on the internal mechanisms by which LMs encode and manipulate representations of mental states to solve (or fail to solve) such tasks (Hu et al., 2025; Gweon et al., 2023).
@@
In this work, we investigate how LMs represent and update characters’ beliefs, which is a fundamental element of ToM (Dennett, 1981; Wimmer & Perner, 1983). For instance, the Sally-Anne test (Baron-Cohen et al., 1985), a canonical test of ToM in humans, evaluates this ability by asking individuals to track Sally’s belief, which diverges from reality due to missing information, and Anne’s belief, which updates based on new observations.
@@
We construct CausalToM, a dataset of simple stories involving two characters, each interacting with an object to change its state, with the possibility of observing one another. We then analyze the internal mechanisms that enable Llama-3-70B-Instruct (Grattafiori et al., 2024) to reason about and answer questions regarding the characters’ beliefs about the state of each object (for a sample story, see Section 3 and for the full prompt refer to Appendix A).
@@
Figure 1: The lookback mechanism is used to perform conditional reasoning. The source token contains information that is copied into two instances via attention to create a pointer and an address. Alongside the address in the residual stream is a payload of information. When necessary, the model retrieves the payload by dereferencing the pointer. Solid lines represent movement of information, while the dotted line indicates the attention “looking back” from pointer to address.
@@
We discover a pervasive computation that performs multiple subtasks, which we refer to as the lookback mechanism. This mechanism enables the model to recall important information only when it becomes necessary. In a lookback, two copies of a single piece of information are transferred to two distinct tokens. This allows attention heads at the latter token to look back at the earlier one when needed and retrieve vital information stored there, rather than transferring that information directly (see Fig. 1).
@@
We identify three key lookback mechanisms that collectively perform belief tracking: 1) Binding lookback (Fig. 2(a)(i)): First the LM assigns ordering IDs (OIs) (Dai et al., 2024) that encode whether a character, object, or state token appears first or second. Then, the character and object OIs are copied to low-rank subspaces of the corresponding state token and the final token residual stream. Later, when the LM needs to answer a question about a character’s beliefs, it uses this information to retrieve the answer state OI. 2) Answer lookback (Fig. 2(a)(ii)): Uses the answer state OI from the binding lookback to retrieve the answer state token value. 3) Visibility lookback (Fig. 6): When an explicit visibility condition between characters is mentioned, the model employs additional reference information called the visibility ID to retrieve information about the observed character, augmenting the observing character’s awareness.
@@
Overall, this work not only advances our understanding of the internal computations in LMs that enable ToM capability but also uncovers a pervasive mechanism that serves as the foundation for executing complex logical reasoning with conditionals. All code and data supporting this study are available at https://belief.baulab.info.
@@ 2 The Lookback Mechanism
((a))
@@
Refer to captionCounterfactualOriginal
((b))
Figure 2: Belief Tracking in Language Models: We task the LM with tracking the beliefs of two characters that manipulate the states of two objects. We hypothesize that the LM solves this task by implementing a causal model with two lookback mechanisms (2(a)). To support our hypothesis, we conduct a causal analysis where we measure whether interventions on a high-level causal model produce the same output as equivalent interventions on the LM. For instance in 2(b), we show results for an experiment distinguishing the pointer and payload in the answer lookback.
(a) Belief Tracking with No Visibility between Characters: Our hypothesized causal model for this kind of story has two lookbacks that operate on ordering IDs (OIs) that encode whether a token appears first or second. In the binding lookback (i), the LM first represents the two events in the story by binding together each character-object-state triple in the residual stream of the state token. When questioned about a particular character and object, the LM looks back to the corresponding triple and retrieves an OI to that state token. Notice that in this lookback, that payload is later used as a pointer, i.e., what a C programmer would call a double pointer. In the answer lookback (ii), the LM dereferences the pointer to the answer token to generate the correct answer. Color indicates the information content, while shape indicates the role of that information in lookback (see Fig. 1), e.g., the state OI is a payload (Refer to caption) in the binding lookback and a pointer/address (Refer to caption) in the answer lookback.
@@
(b) Answer Lookback Pointer and Payload: To test our hypothesized causal model, we run the LM on pairs of slightly different stories, and then intervene by patching a specific representation state from the counterfactual run to the original run, observing any change in the output. The causal model predicts that if we alter the “answer payload Refer to caption” of the original to instead take the value of the counterfactual answer payload, the output should change from coffee to tea; the gray curve in the line plot shows this does occur with $p\approx 1.0$ when patching states at the “:” token beyond layer 56, providing evidence that the answer payload resides in those states. On the other hand the causal model predicts that taking the counterfactual “answer pointer Refer to caption” would change the original run output from coffee to beer—a new output that matches neither the original nor the counterfactual!—and we do see this surprising effect, again with $p\approx 1.0$, when patching layers between 34 and 52, providing strong evidence that the answer pointer is encoded at those layers. Collected over $N=80$ samples, these measurements suggest the Answer Lookback occurs between layers 52 and 56. Furthermore the representations of the causal variables are small: the interventions can be localized even further to subspaces of dimension 33 (payload) or 20 (pointer), tiny portions of the 8192-dimensional state space.
Our investigation of belief tracking uncovers a recurring pattern of computation that we call the lookback mechanism. [^1] Although this mechanism may resemble induction heads (Elhage et al., 2021; Olsson et al., 2022), they differ fundamentally. In induction heads, information from a previous token occurrence is passed only to the subsequent token through, without being duplicated to its next occurrence. In contrast, the lookback mechanism copies the same information not only to the location where the vital information resides but also to the target location that needs to retrieve that information. Here we give a brief overview of this mechanism; subsequent sections provide detailed experiments and analyses. In lookback, source information is copied (via attention) into an address copy in the residual stream of a recalled token and a pointer copy in the residual stream of a lookback token that occurs later in the text. The LM places the address alongside a payload of the recalled token’s residual stream that can be brought forward to the lookback token if necessary. Fig. 1 schematically describes a generic lookback.
That is, the LM can use attention to dereference the pointer and retrieve the payload present in the residual stream of the recalled token (that might contain aggregated information from previous tokens), bringing it to the residual stream of the lookback token. Specifically, the pointer at the lookback token forms an attention query vector, while the address at the recalled token forms a key vector. Because the pointer and the address are copies of the same source information, they would have a high dot-product, hence a QK-circuit (Elhage et al., 2021) is established forming a bridge from the lookback token to the recalled token. The LM uses this bridge to move the payload that contains information needed to complete the subtask through the OV-circuit.
To develop an intuition for why an LM would learn to implement lookback mechanisms to solve reasoning tasks such as our belief tracking task, consider that during training LMs process text in sequence with no foreknowledge of what might come next. Then, it would be useful to mark addresses alongside payloads that might be useful for downstream tasks. In our setting, the LM constructs a representation of the story without any knowledge of what questions it may be asked about, so the LM concentrates pieces of information in the residual stream of certain tokens which later become payloads and addresses. When the question text is reached, pointers are constructed that reference this crucial story information and dereference it as the answer to the question.
3 Preliminaries
Dataset    Existing datasets for evaluating ToM capabilities of LMs are designed for behavioral testing and lack counterfactual pairs needed for causal analysis (Kim & Sundar, 2012). To address this, we constructed CausalToM, a structured dataset of simple stories, where each story involves two characters, each interacting with a distinct object causing the object to take a unique state. For example: “Character1 and Character2 are working in a busy restaurant. To complete an order, Character1 grabs an opaque Object1 and fills it with State1. Then Character2 grabs another opaque Object2 and fills it with State2.” We then ask the LM to reason about one of the characters’ beliefs regarding the state of an object: “What does Character1 believe Object2 contains?” We analyze the LM’s ability to track characters’ beliefs in two distinct settings. (1) No Visibility, where both characters are unaware of each other’s actions, and (2) Explicit Visibility where explicit information about whether a character can/cannot observe the other’s actions is provided, e.g., “Bob can observe Carla’s actions. Carla cannot observe Bob’s actions.” We also provide general task instructions (e.g., answer unknown when a character is unaware); refer to Appendix A & B for the full prompt and additional dataset details. Our experiments analyze the Llama-3-70B-Instruct model in half-precision, using NNsight (Fiotto-Kaufman et al., 2025). The model demonstrates a high behavioral performance on both the no-visibility and explicit-visibility settings, achieving accuracy of 95.7% and 99% respectively. For all subsequent experiments, we filter out samples that the model fails to answer correctly.
Causal Mediation Analysis
Our goal is to develop a mechanistic understanding of how Llama-3-70B-Instruct reasons about characters’ beliefs and answers related questions (Saphra & Wiegreffe, 2024). A key method for conducting causal analysis is interchange interventions (Vig et al., 2020; Geiger et al., 2020; Finlayson et al., 2021), in which the LM is run on paired examples: an original input $\mathbf{o}$ and a counterfactual input $\mathbf{c}$ and certain internal activations in the LM run on the original are replaced with those computed from the counterfactual.
Drawing inspiration from existing literature (Vig et al., 2020; Meng et al., 2022; Wang et al., 2023), we begin our analysis by performing interchange interventions with counterfactuals that are identical to the original except for key input tokens. We trace the causal path from these key tokens to the final output. This is a type of Causal Mediation Analysis (Mueller et al., 2024). Specifically, we construct a counterfactual dataset where $\mathbf{o}$ contains a question about the belief of a character not mentioned in the story, while $\mathbf{c}$ is identical except that the story includes the queried character. The expected outcome of this intervention is a change in the final output of $\mathbf{o}$ from unknown to a state token, such as beer. We conduct similar interchange interventions for object and state tokens (refer to Appendix C for more details).
Figure 3 presents the aggregated results of this experiment for the key input tokens Character1, Object1, and State1. The cells are color-coded to indicate the interchange intervention accuracy (IIA; Geiger et al., 2022). Even at this coarse level of analysis, several significant insights emerge: 1) Information from the correct state token (beer) flows directly from its residual stream to that of the final token in later layers, consistent with prior findings  (Lieberum et al., 2023; Prakash et al., 2024); 2) Information associated with the query character and the query object is retrieved from their earlier occurrences and passed to the final token before being replaced by the correct state token.
Desiderata Based Patching via Causal Abstraction
The causal mediation experiments provide a coarse-grained analysis of how information flows from an input token to the output but do not identify what that information is. A fact about transformers is that the input to the first layer contains input tokens and the output from the final layer contains the output token, but what is the information content of representations along the active causal path in between the input and output?
To answer this question, we turn to Causal Abstraction (Geiger et al., 2021; 2024). We align the variables of a high-level causal model with the LM’s internal activations and verify the alignment by conducting targeted interchange interventions for each variable. Specifically, we perform aligned interchange interventions at both levels: interventions that target high-level causal variables and interventions that modify low-level features of the LM’s hidden activations. If the LM produces the same output as the high-level causal model under these aligned interventions, it provides evidence supporting the hypothesized causal model. The effect of these interventions is quantified using IIA, which measures the proportion of instances where intervened high-level causal model and low-level LM have the same output (refer to Appendix D for more details about the causal abstraction framework and Appendix E for the belief tracking causal model).
In addition to performing interchange interventions on entire residual stream vectors in LMs, we also intervene on specific subspaces to further localize causal variables. To identify the subspace encoding a particular variable, we employ the Desiderata-based Component Masking (De Cao et al., 2020; Davies et al., 2023; Prakash et al., 2024) technique, which learns a sparse binary mask over the internal activation space by maximizing the logit of the causal model output token. Specifically, we train a mask to select the singular vectors of the activation space that encode a high-level variable (see Appendix F for details).
4 Belief Tracking via Ordering IDs and Lookback Mechanisms
The LM solves the no visibility setting of the belief tracking task using three key mechanisms: Ordering ID assignment, binding lookback, and answer lookback. Figure 2(a) illustrates the hypothesized high-level causal model implemented by the LM, which we evaluate in the following subsections. The LM first assigns ordering IDs (OIs; [Uncaptioned image], [Uncaptioned image], [Uncaptioned image]) to each character, object, and state in the story that encode their order of appearance (e.g., the second character Carla is assigned [Uncaptioned image]). These OIs are used in two lookback mechanisms. (i) Binding lookback: Address copies of each character OI ([Uncaptioned image]) and object OI ([Uncaptioned image]) are placed alongside their corresponding state OI payload ([Uncaptioned image]) in the residual stream of each state token, binding together each character-object-state triple. When the model is asked about the belief of a specific character about a specific object, it moves pointer copies of the corresponding OIs ([Uncaptioned image], [Uncaptioned image]) to the final token’s residual stream. These pointers are dereferenced, bringing the correct state OI into the final token residual stream. (ii) Answer lookback: An address copy of the state OI ([Uncaptioned image]) is alongside the state token payload ([Uncaptioned image]) in the residual stream of the correct state token, while a pointer copy ([Uncaptioned image]) is moved to the final token residual stream via the binding lookback. The pointer is dereferenced, bringing the answer state token payload into the final token residual stream, which is predicted as the final output.
Refer to Appendix E for pseudocode defining the causal model for the belief tracking task. In Appendices L and K, we show that parts of our analysis generalizes to the Llama-3.1-405B-Instruct model and the BigToM dataset (Gandhi et al., 2024).
4.1 Ordering ID Assignment
The LM assigns an Ordering ID (OI; Dai et al. 2024) to the character, object, and state tokens. These OIs, encoded in a low-rank subspace of the internal activation, serve as a reference that indicates whether an entity is the first or second of its type independent of its token value. For example, in Fig. 2(a), Bob is assigned the first character OI, while Carla receives the second. In the subsequent subsections and Appendices G & H, we validate the presence of OIs through multiple experiments, where intervening on tokens with identical token values but different OIs alters the model’s internal computation, leading to systematic changes in the final output predicted by our high-level causal model. The LM then uses these OIs as building blocks, feeding them into lookback mechanisms to track and retrieve beliefs.
4.2 Uncovering the Binding Lookback Mechanism

### v2

@@ Abstract
How do language models (LMs) represent characters’ beliefs, especially when those beliefs may differ from reality? This question lies at the heart of understanding the Theory of Mind (ToM) capabilities of LMs. We analyze LMs’ ability to reason about characters’ beliefs using causal mediation and abstraction. We construct a dataset, CausalToM, consisting of simple stories where two characters independently change the state of two objects, potentially unaware of each other’s actions. Our investigation uncovers a pervasive algorithmic pattern that we call a lookback mechanism, which enables the LM to recall important information when it becomes necessary. The LM binds each character-object-state triple together by co-locating their reference information, represented as Ordering IDs (OIs), in low-rank subspaces of the state token’s residual stream. When asked about a character’s beliefs regarding the state of an object, the binding lookback retrieves the correct state OI and then the answer lookback retrieves the corresponding state token. When we introduce text specifying that one character is (not) visible to the other, we find that the LM first generates a visibility ID encoding the relation between the observing and the observed character OIs. In a visibility lookback, this ID is used to retrieve information about the observed character and update the observing character’s beliefs. Our work provides insights into belief tracking mechanisms, taking a step toward reverse-engineering ToM reasoning in LMs.
@@ 1 Introduction
Theory of Mind (ToM), the ability to infer others’ mental states, is an essential aspect of social and collective intelligence (Premack & Woodruff, 1978; Riedl et al., 2021). Recent studies have established that LMs can solve some tasks requiring ToM reasoning (Street et al., 2024; Strachan et al., 2024a; Kosinski, 2024), while others have highlighted shortcomings (Ullman, 2023; Sclar et al., 2025; Shapira et al., 2024, inter alia). Previous studies primarily rely on behavioral evaluations, which do not shed light on the internal mechanisms by which LMs encode and manipulate representations of mental states to solve (or fail to solve) such tasks (Hu et al., 2025; Gweon et al., 2023).
@@
In this work, we examine how LMs internally represent and track beliefs of characters, a core aspect of ToM (Dennett, 1981; Wimmer & Perner, 1983). A classic example is the Sally-Anne test (Baron-Cohen et al., 1985), which evaluates ToM in humans by assessing whether individuals can track conflicting beliefs: Sally’s belief, which diverges from reality because of missing information, and Anne’s belief, which is updated based on new observations. Our goal is to determine whether LMs learn a systematic solution to such tasks or rely on superficial statistical association.
@@
We construct CausalToM, a dataset of simple stories involving two characters, each interacting with an object to change its state, with the possibility of observing one another. We then analyze the internal mechanisms that enable Llama-3-70B-Instruct and Llama-3.1-405B-Instruct (Grattafiori et al., 2024) to reason about and answer questions regarding the characters’ beliefs about the state of each object (for a sample story, see Section 3 and for the full prompt refer to Appendix A).
@@
Figure 1: The lookback mechanism performs conditional reasoning; The source token contains reference information that is copied into two instances, creating a pointer and an address. Next to the address in the residual stream is a payload. When necessary, the model retrieves the payload by dereferencing the pointer. Solid lines represent information flow, while the dotted line indicates the attention “looking back” from pointer to address.
@@
Our findings provide strong evidence for a systematic solution to belief tracking. We discover that LMs use a pervasive computation, which we refer to as the lookback mechanism, for belief tracking. This mechanism enables the model to recall important information at a later stage. In a lookback, two copies of a single piece of information are transferred to two distinct tokens. This allows attention heads at the latter token to look back at the earlier one when needed and retrieve vital information stored there, rather than transferring it directly (see Fig. 1).
@@
We identify three key lookback mechanisms that collectively perform belief tracking: 1) Binding lookback (Fig. 3(i)): First, the LM assigns ordering IDs (OIs; Dai et al. 2024) that encode whether a character, object, or state token appears first or second. Then, the character and object OIs are copied to the corresponding state token and the final token residual stream. Later, when the LM needs to answer a question about a character’s beliefs, it uses this information to retrieve the answer state OI. 2) Answer lookback (Fig. 3(ii)): Uses the answer state OI from the binding lookback to retrieve the answer state token value. 3) Visibility lookback (Fig. 7): When a visibility condition between characters is mentioned, the model employs additional reference information called the visibility ID to retrieve information about the observed character, augmenting the observing character’s awareness.
@@
Overall, this work not only advances our understanding of the internal computations in LMs that enable ToM but also uncovers a pervasive mechanism that plays a foundational role for in-context reasoning. All code and data supporting this study are available at https://belief.baulab.info.
@@ 2 The Lookback Mechanism
Our investigation uncovers a recurring pattern of computation that we call the lookback mechanism. [^1] Although this mechanism may resemble induction heads (Elhage et al., 2021; Olsson et al., 2022), it differs fundamentally. In induction heads, information from a previous token occurrence is passed only to the subsequent token, without being duplicated to its next occurrence. In contrast, the lookback mechanism copies the same information not only to the location where the vital information resides but also to the target location that needs to retrieve that information. In lookback, a source reference is copied (via attention) into an address copy in the residual stream of the recalled token and a pointer copy in the residual stream of the lookback token that occurs later in the text. The LM places the address alongside a payload in the recalled token’s residual stream that can be brought forward to the lookback token if necessary. Fig. 1 shows a generic lookback.
That is, the LM can use attention to dereference the pointer and retrieve the payload present in the residual stream of the recalled token (which might contain aggregated information from previous tokens), bringing it to the residual stream of the lookback token. Specifically, the pointer at the lookback token forms an attention query vector, while the address at the recalled token forms a key vector. The pointer and address are not necessarily exact copies of the source reference, but they do have a high dot product after being transformed by a query or key attention matrix, respectively. Hence, a QK-circuit (Elhage et al., 2021) is established, forming a bridge from the lookback token to the recalled token. The LM uses this bridge to move the payload that contains information needed to complete the subtask through the OV-circuit.
To develop an intuition for why an LM would learn to implement lookback mechanisms, consider that during training, LMs process text in sequence with no foreknowledge of what might come next. Instead of trying to resolve every possible future question about the current context, it would be useful to place addresses alongside payloads that might be useful to remember in the future when performing a variety of downstream tasks. In our setting, the LM constructs a representation of a story without any certainty about the questions it may later be asked about that story, so the LM localizes pivotal information in the residual stream of certain tokens, which later become payloads and addresses. When the question text is reached, pointers are constructed that reference this crucial story information and dereference it to find an answer to the question.
3 Experimental Setup: Dataset, Models, and Methods
Dataset    Existing datasets for evaluating ToM capabilities of LMs are designed for behavioral testing and lack counterfactual pairs needed for causal analysis (Kim & Sundar, 2012). To address this problem, we construct CausalToM, a structured dataset of simple stories, where each story involves two characters, each interacting with a distinct object causing the object to take a unique state. For example: “Character1 and Character2 are working in a busy restaurant. To complete an order, Character1 grabs an opaque Object1 and fills it with State1. Then Character2 grabs another opaque Object2 and fills it with State2.” We then ask the LM to reason about one of the characters’ beliefs regarding the state of an object: “What does Character1 believe Object2 contains?” We analyze the LM’s ability to track characters’ beliefs in two distinct settings. (1) No Visibility, where both characters are unaware of each other’s actions, and (2) Explicit Visibility, where explicit information about whether a character can/cannot observe the other’s actions is provided, e.g., “Bob can observe Carla’s actions. Carla cannot observe Bob’s actions.” We also provide general task instructions (e.g., answer unknown when a character is unaware); refer to Appendices A & B for the full prompt and additional dataset details. All subsequent experiments are conducted on 80 samples that the model answers correctly. We also demonstrate generalization of the mechanism to BigToM dataset (Gandhi et al., 2024) in Appendix K.
Models    Our experiments analyze Llama-3-70B-Instruct and Llama-3.1-405B-Instruct models in FP16 and INT8 precision, respectively, using NNsight (Fiotto-Kaufman et al., 2025). Results for Llama-3.1-405B-Instruct can be found in Appendix L. Both models demonstrate strong behavioral performance in the no-visibility and explicit-visibility settings. We do not examine smaller models, as they are unable to coherently solve the CausalToM task.
Causal Mediation Analysis    Our goal is to develop a mechanistic understanding of how LMs reason about characters’ beliefs and answer related questions (Saphra & Wiegreffe, 2024). A key method for conducting causal analysis is interchange interventions (Vig et al., 2020; Geiger et al., 2020; Finlayson et al., 2021), in which the LM is run on paired examples: an original input $\mathbf{o}$ and a counterfactual input $\mathbf{c}$, and certain internal activations in the LM run on the original input are replaced with those computed from the counterfactual, a process also known as activation patching. We begin our analysis by tracing information flow from key input tokens to the final output, by performing interchange interventions on the residual vectors. Specifically, we construct an intervention dataset where $\mathbf{o}$ contains a question about the belief of a character not mentioned in the story, while the story in $\mathbf{c}$ includes the same queried character, as shown in Fig. 2. The expected outcome of this intervention is a change in the final output of $\mathbf{o}$ from unknown to a state token, such as beer. We conduct similar interchange interventions for object and state tokens (refer to Appendix C for details).
Figure 2 presents the aggregated results of this experiment for the key input tokens Character1, Object1, and State1. The cells are color-coded to indicate the interchange intervention accuracy (IIA; Geiger et al., 2022). Even at this coarse level of Causal Mediation Analysis (Mueller et al., 2024; Vig et al., 2020; Meng et al., 2022), several significant insights emerge: 1) Information from the correct state token (beer) flows directly from its residual stream to that of the final token in later layers, consistent with prior findings  (Lieberum et al., 2023; Prakash et al., 2024); 2) Information associated with the query character and the query object is retrieved from their earlier occurrences and passed to the final token before being replaced by the correct state token.
Desiderata Based Patching via Causal Abstraction    The causal mediation experiments provide a coarse-grained analysis of where information flows, but do not identify what information is being transferred. In a transformer, the first layer represents the input and the last layer represents the output, but we wish to know: what is represented in the middle? We analyze the internal mechanism using Causal Abstraction (Geiger et al., 2021; 2024); First, we hypothesize a high-level causal model of the computational steps from input to output (Sec. 4), and then align its variables with the LM’s internal activations (Sec. 5). We test the alignment through targeted interchange interventions on causal variables in the hypothesized model and hidden activations in the LM. If the LM produces the same output as the causal model under these aligned interventions, it provides evidence supporting the hypothesized causal model. We quantify this effect using interchange intervention accuracy (IIA; Geiger et al., 2022), which measures the proportion of cases where the intervened causal model and intervened LM agree. See Appendix D for more details.
In addition to measuring IIA on entire residual stream vectors, we also intervene on localized subspaces to further isolate causal variables. To identify the subspace of a specific variable, we employ Desiderata-based Component Masking (De Cao et al., 2020; Davies et al., 2023; Prakash et al., 2024). This method learns a sparse binary mask over the activation space that maximizes the logit of the hypothesized causal model output. We train a mask to select singular vectors of the activation space that encode a high-level variable (see Appendix F for details). Our experiments in Sec. 5 report both interventions on the full residual stream and on the identified subspaces.
4 Hypothesized High-Level Causal Model of Belief Tracking
Figure 3: Belief Tracking with no visibility between characters. We hypothesize that the LM tracks beliefs using two lookback mechanisms. First, in (i) Binding lookback, LM binds together each character-object-state triple in the state token residual stream. When asked about a specific character-object pair, the LM looks back to the corresponding OIs to retrieve the correct state OI. Second, in (ii) Answer lookback, LM dereferences that state OI (used as a pointer) to retrieve the token value of the correct state. Colors indicate information type, shapes indicate role of information in lookback (see Fig. 1), e.g., state OI is a payload (Refer to caption) in (i) and a pointer-address (Refer to caption) in (ii).
Here we start with an overview of our hypothesized causal model of belief-tracking when characters are not aware of each other’s actions. The causal model is an algorithmic process that has variables with structural roles that do not refer to the details of a transformer architecture. Appendix E presents the full pseudocode of the causal model. In Section 5, we will present experiments to verify that the causal model’s variables align with representations in the transformer.
Belief tracking begins when the causal model assigns ordering IDs (OIs; [Uncaptioned image], [Uncaptioned image], [Uncaptioned image]) to each character, object, and state token, marking their order of appearance. For instance, in the example in Fig. 3, Bob is assigned first character OI ([Uncaptioned image]), and Carla is assigned the second character OI([Uncaptioned image]). Then it uses these OIs in two lookback mechanisms:
(i) Binding lookback. The causal model creates address copies of each character OI ([Uncaptioned image]) and object OI ([Uncaptioned image]) that are bound to the state OI (Binding Payload, [Uncaptioned image]), creating a character–object–state triple. When a question is asked about a character and object, the causal model creates pointer copies of that character and object OIs (Binding Pointers [Uncaptioned image], [Uncaptioned image]) and dereferences them to retrieve the state OI.
(ii) Answer lookback. The causal model creates an address copy of the state OI (Answer Address [Uncaptioned image]) that is bound to the state token (Answer Payload, [Uncaptioned image]). Through the binding lookback, a pointer copy of this OI ([Uncaptioned image]) is created. The causal model dereferences the pointer to retrieves the correct state token payload as the final output.
5 Verifying the Hypothesized Causal Model of Belief Tracking
We test our hypothesized causal model by localizing its variables within the transformer’s neural representations. Specifically, we localize the addresses, pointers, and payloads of the (i) binding lookback and (ii) answer lookbacks within the LM’s internal activations. In Fig. 3, we show a trace of the causal model run on an input overlayed onto a schematic of a transformer architecture. This visualizes the alignment between variables in the causal model and locations in the LM residual stream that the experiments in the remaining of this paper will support. In the binding lookback, the character and object OI addresses are realized in the residual stream of the state token. The pointer copies are brought forward to the last token residual stream where they are dereferenced via attention to bring forward the correct state OI payload. In the answer lookback, the address copy of the state OI is in the state token residual stream while the pointer copy is in the last token residual stream.
Each of the following experiments localizes the presence of specific ordering IDs (OIs) and verifies their roles as hypothesized by our causal model. We do this by targeted interchange intervention experiments on the causal model and the LM. We copy hidden states between identical tokens (for example, replacing the representation of “:” in one context with the representation of “:” in another context, as in Fig. 4). When this intervention causes the LM’s have the same output as the causal model under an interchange intervention on OI variables, we have evidence that the OI is carrying out the hypothesized role. Each experiment reports the effects of $n=80$ different cases with the same structure, and the effect is measured at every layer.
Because the last step of the causal model is easiest to understand, we proceed through the experiments in reverse order, beginning with an experiment to verify the final “answer lookback” stage. After this instructive starting point, we work backward to verify the earlier steps of the model. Additional results can be found in Appendix G and H.
5.1 Step ii: Answer Lookback – Retrieving the Correct State
@@
Refer to caption(a) Intervention Input Example(b) Layer-wise Intervention ResultsCounterfactualOriginal
@@
Figure 4: Answer Lookback Pointer and Payload: The causal model predicts that if we alter the “Answer Payload Refer to caption” of the original to instead take the value of the counterfactual answer payload, the output should change from coffee to tea; the gray curve in the line plot shows this does occur when patching residual vectors at the “:” token beyond layer $56$, providing evidence that the answer payload resides in those states. On the other hand the causal model predicts that taking the counterfactual “Answer Pointer Refer to caption” would change the original run output from coffee to beer—a new output that matches neither the original nor the counterfactual!—and we do see this surprising effect, again when patching layers between $34$ and $52$, providing strong evidence that the answer pointer is encoded at those layers. These results suggest the Answer Lookback occurs between layers 52 and 56.
Localizing the Answer Payload    We first verify the presence of the correct Answer Payload [Uncaptioned image] at the deepest layer representation of final token “:”. To do so, we run an interchange intervention experiment shown in Fig. 4a in which the counterfactual example $\mathbf{c}$ swaps the order of the characters and objects of the original example $\mathbf{o}$ and also replaces the state (drinks) tokens with new values. If the Answer Payload is correctly localized, swapping it should cause the answer of the counterfactual (e.g., tea) to replace the answer of the original example (e.g., coffee). The gray line in Fig. 4b shows that this output change is observed in every one of $n=80$ cases, both when intervening on the full residual stream and on the identified subspace. However, not at every layer: the information is only present after layer $56$, indicating that before this stage, the transformer has not yet retrieved the correct answer payload into the residual stream. That is consistent with our hypothesis that at early steps, the OI has not yet been dereferenced. At an earlier stage, we expect to see an Answer Pointer.
Localizing the Pointer Information    To identify the Answer Pointer [Uncaptioned image] before it is dereferenced to bring the payload (state token value), we examine the representations of “:” at layers earlier than $56$. Our causal model provides the hypothesis: if the Answer Pointer is present, then patching the pointer from the counterfactual run into the original run should redirect the LM to attend to the location of the correct counterfactual state and fetch its payload. For example, in Fig. 4a the counterfactual pointer references the first presented state. When we patch it into the original story, we expect the model’s answer to change to beer rather than coffee. The colored line in Fig. 4b confirms that this effect is consistently observed when patching any layer between 345234-52 (both when patching the full residual stream and the identified subspace), supporting our hypothesis that these layers encode the Answer Pointer information at the final token, rather than directly transferring token values.
5.2 Step i: Binding Lookback – Linking Characters, Objects, and States
Localizing the Address and Payload    In this experiment, we verify the presence of the address copies of the character and object OIs  as well as the payload (state OI) at the state token residual stream (recalled token, Fig. 3). As illustrated in Fig. 5a, we construct an intervention dataset where each example consists of an original input $\mathbf{o}$ with an answer that is not unknown and a counterfactual input $\mathbf{c}$ where the character, object, and state token values are identical, except the ordering of the two story sentences is swapped while the question remains unchanged. The expected LM’s output predicted by our hypothesized causal model is the other state token in the original example, e.g., beer. That is because patching the address and payload values at each state token, without changing the pointer, makes the LM dereference the other state token. As a result, the model’s output should flip to the other state token in the original input.